# The externally definable Ramsey property and fixed points on type spaces

### v3

@@ 3.2. Simplifications via quantifier elimination
Let $M$ be an ultrahomogeneous structure. Suppose $\operatorname{Th}(M)$ has QE. Let $\bar{a}\in\operatorname{Age}(M)$. Suppose $M$ has $\bar{a}$-Ramsey for externally atomic-definable $2$-colourings. Then M has $\bar{a}$-Ramsey for all externally definable $2$-colourings
@@
Apply Lemma 3.8 to the set of quantifier-free formulae. ∎
@@
Assume $M$ is ultrahomogeneous. We consider the case of equality. The $2$-colourings externally defined by the atomic formula $\phi(\bar{x},\bar{y})\equiv x_{0}=x_{1}$ or by $\phi(\bar{x},\bar{y})\equiv y_{0}=y_{1}$ are constant, so immediately have $\bar{a}$-Ramsey. However, it is possible for $2$-colourings externally defined by the atomic formula $\phi(\bar{x},\bar{y})\equiv x=y$ to not have point-Ramsey. For example, let $\mathcal{L}$ consist of a single unary predicate $R$, which we think of as the “red points” of a set, and let $M$ be the Fraïssé limit of the class of finite $\mathcal{L}$-structures $D$ with $|R^{D}|\leq 2$, i.e. with at most two red points. Let $a\in M$ be a red point, and let $B\in\operatorname{Age}(M)$ consist of two red points. Then the $2$-colouring of $\binom{M}{a}$ externally defined by $a\in M$ and the formula $\phi(x,y)\equiv x=y$ has no monochromatic copy of $B$.
However, if $\operatorname{Age}(M)$ has the strong JEP, then $2$-colourings externally defined by $e\in M$ and $x=y$ have point-Ramsey: any copy of $B$ which does not contain $e$ is monochromatic. Thus, if $\operatorname{Age}(M)$ has the strong JEP, all atomic formulae involving equality automatically have $\bar{a}$-Ramsey.
@@
Let $M$ be an ultrahomogeneous $\mathcal{L}$-structure, where $\mathcal{L}$ has maximum arity $t\geq 2$. Suppose $\operatorname{Th}(M)$ has QE. Suppose $M$ has externally definable $\bar{a}$-Ramsey for all $\bar{a}\in\operatorname{Age}(M)$ with $|\bar{a}|<t$. Then $M$ has the externally definable Ramsey property.
@@
By Lemma 3.10, it suffices to consider externally atomic-definable colourings. Let $\bar{a}\in\operatorname{Age}(M)$, and let $\chi:\binom{M}{\bar{a}}\to 2$ be externally defined by some $\bar{e}\in N\succeq M$ and an atomic formula $\phi(\bar{x},\bar{y})$. Take subsequences $(x_{r(i)}),(y_{s(i)})$ of $\bar{x},\bar{y}$ respectively, consisting of the free variables that actually occur in $\phi(\bar{x},\bar{y})$. Let $u$ be the length of the subsequence $(x_{r(i)})$, and let $v$ be the length of the subsequence $(y_{s(i)})$.
@@
Let $M$ be an ultrahomogeneous $\mathcal{L}$-structure, where $\mathcal{L}$ has maximum arity $t\geq 2$. Suppose $\operatorname{Th}(M)$ has QE. Suppose, for $\bar{a}\in\operatorname{Age}(M)$ with $|\bar{a}|<t$, that $M$ has $\bar{a}$-Ramsey for externally atomic-definable $2$-colourings. Then $M$ has the externally definable Ramsey property.
(This resembles [16, Proposition 3.17].)
@@ 4. The fixed points on type spaces property (FPT)
Let $M$ be an $L$-structure, and let $G=\operatorname{Aut}(M)$ with the pointwise convergence topology. Then $G$ acts continuously on the Stone spaces $S_{n}(M)$, with the action given by
@@
Note that we define the action of $G$ on $L(M)$-formulae as $g\cdot\phi(\bar{x},\bar{m})=\phi(\bar{x},g\bar{m})$.
@@
Let $M$ be an $\mathcal{L}$-structure with automorphism group $G=\operatorname{Aut}(M)$. We say that $G$ has the fixed points on type spaces property (FPT) if every subflow of $G\curvearrowright S_{n}(M)$, $n\geq 1$, has a fixed point.
We also define a more restricted version of this property for particular $n$. Let $n\geq 1$. We say that $G$ has FPTn if every subflow of $G\curvearrowright S_{n}(M)$ has a fixed point.
@@ 5. FPT and the externally definable Ramsey property
    if $\operatorname{Aut}(M)$ has FPT, then $M$ has the externally definable Ramsey property;
@@
    if $\operatorname{Age}(M)$ is countable, $M$ has the externally definable Ramsey property iff $\operatorname{Aut}(M)$ has FPT.
@@
(1): Let $\bar{a}\in\operatorname{Age}(M)$. We may assume $\bar{a}\subseteq M$. By Lemma 3.7, it suffices to show $\bar{a}$-Ramsey for externally definable $2$-colourings. Let $\chi:\binom{M}{\bar{a}}\to 2$ be externally defined by $e\in N\succeq M$ and $\phi(\bar{x},\bar{y})$, where $\phi(\bar{x},\bar{y})$ has free variables among $\bar{x},\bar{y}$, but including all of $\bar{x}$. We show that for $B\in\operatorname{Age}(M)$, there exists a $\chi$-monochromatic copy $B^{\prime}\subseteq M$ of $B$.
Let $p(\bar{x})=\operatorname{tp}_{N}(\bar{e}/M)$. As $\operatorname{Aut}(M)$ has FPT, there is a $G$-invariant type $q(\bar{x})\in\overline{Gp(\bar{x})}$. Thus:
@@
(2): Take $p(\bar{x})\in S_{n}(M)$. We will show that there is a $G$-invariant $q(\bar{x})\in\overline{Gp(\bar{x})}$.
Let $N\succeq M$ and $\bar{e}\in N$ be such that $\operatorname{tp}_{N}(\bar{e}/M)=p(\bar{x})$. For $\bar{a}\in\operatorname{Age}(M)$ and for $\phi(\bar{x},\bar{y})$ an $\mathcal{L}$-formula with $|\bar{x}|=|\bar{e}|$ and $|\bar{y}|=|\bar{a}|$, let $\chi_{\phi,\bar{a}}:\binom{M}{\bar{a}}\to 2$ denote the colouring externally defined by $\bar{e}\in N\succeq M$ and $\phi(\bar{x},\bar{y})$. Here we assume that we have fixed the interpretation of $\bar{x}$ as $\bar{e}$: i.e. for any $\mathcal{L}$-formula $\psi(\bar{z})$, we partition $\bar{z}$ into $\bar{z}\cap\bar{x}$ and $\bar{y}=\bar{z}\setminus\bar{x}$, and thereby consider $\psi(\bar{z})$ as a formula $\psi(\bar{x},\bar{y})$. Note that for $\bar{a},\bar{b}\in\operatorname{Age}(M)$ with $\bar{a}\cong\bar{b}$, $\chi_{\phi,\bar{a}}=\chi_{\phi,\bar{b}}$.
@@
Let ${(\lambda_{r})}_{r<\omega}$ be given by the above Claim 2. Define $q(\bar{x})$ as follows. For $r<\omega$, let $\psi_{r}(\bar{x},\bar{y})=\begin{cases}\phi_{r}(\bar{x},\bar{y})&\lambda_{r}=0\\ \neg\phi_{r}(\bar{x},\bar{y})&\lambda_{r}=1\end{cases}$  and add $\psi_{r}(\bar{x},\bar{a}^{\prime}_{r})$ to $q(\bar{x})$ for all $\bar{a}^{\prime}_{r}\in\binom{M}{\bar{a}_{r}}$. By (2), it is immediate that $q(\bar{x})$ is $G$-invariant. To show that $q(\bar{x})$ is a complete type, it suffices to show that $q(\bar{x})\cup\operatorname{Th}_{L(M)}(M)$ is consistent. Take $s,t\in\mathbb{N}$. For $i<s,j<t$, take $\bar{a}_{i,j}\in\binom{M}{\bar{a}_{i}}$. We show that $\{\psi_{i}(\bar{x},\bar{a}_{i,j}):i<s,j<t\}\cup\operatorname{Th}_{L(M)}(M)$ is consistent, and this suffices by compactness. Take finite $B\subseteq M$ with $\bar{a}_{i,j}\subseteq B$ for $i<s,j<t$. Then, by Claim 2, there is $B^{\prime}\subseteq M$ and an isomorphism $f:B\to B^{\prime}$ with $\chi_{i}|_{B^{\prime}}=\{\lambda_{i}\}\text{ for }i<s$. Using ultrahomogeneity of $M$, there is $g\in\operatorname{Aut}(M)$ extending $f$. As $\chi_{i}|_{B^{\prime}}=\{\lambda_{i}\}$ for $i<s$, we have $N\models\psi_{i}(\bar{e},\bar{a}^{\prime}_{i})\text{ for }i<s\text{ and }\bar{a}^{\prime}_{i}\in\binom{B^{\prime}}{\bar{a}_{i}}$, and hence $N\models\psi_{i}(\bar{e},g\bar{a}_{i,j})\text{ for }i<s,j<t$.
@@
We note that, for the case of $\aleph_{0}$-saturated Fraïssé structures, Theorem 5.1 can be deduced from [16, Theorem 3.14] (personal communication, K. Krupiński). The proof of Theorem 5.1 given here uses elementary techniques, and applies to Fraïssé structures generally (or indeed ultrahomogeneous structures with countable age). Several of the examples in the following section are not $\aleph_{0}$-saturated, so [16, Theorem 3.14] cannot be applied, and here Theorem 5.1 becomes relevant. In Section 7.1, we give a general recipe for constructing such examples.
@@ 6. Examples
$\bar{a}$-Ramsey for $|\bar{a}|<t$, where $t$ is the maximum arity of $\mathcal{L}$;
@@
6.1. Finite binary languages: point-Ramsey
@@
Let $\mathcal{L}$ be a language consisting of finitely many binary relation symbols. Let $M$ be a Fraïssé $\mathcal{L}$-structure with point-Ramsey. Then $M$ has the externally definable Ramsey property.
@@
By Lemma 2.6, $\operatorname{Th}(M)$ has QE. Thus, by Proposition 3.13, we need only check point-Ramsey for $2$-colourings of the points of $M$ which are externally defined by atomic formulae, and these form a subset of all $2$-colourings of points of $M$. As $M$ has point-Ramsey, we therefore have the externally definable Ramsey property. ∎
@@
The random graph has point-Ramsey.
@@
Denote the random graph by $M$. We show that $M$ has the pigeonhole property (see Definition 2.17), a stronger property than point-Ramsey. The following is from [4, Proposition 1.3.3] – we include it due to its brevity. Recall that the random graph has the witness property:
@@
One can also quickly prove point-Ramsey for the random graph via the lexicographic product: see [29, Theorem 12.1]).
There are many other Fraïssé structures in finite binary relational languages with point-Ramsey, and therefore the externally definable Ramsey property. For example:
  • the random tournament: this structure has the pigeonhole property by an argument as for the random graph, and we can also prove point-Ramsey via the lexicographic product;
  • the $K_{n}$-free random graph ($n\geq 3$): this follows via Folkman’s theorem, the proof of which adapts the lexicographic product proof for the random graph ([8]; a clear exposition following [15] is [29, Theorem 12.3]); alternatively, we may use the Nešetřil-Rödl theorem as in Section 6.3;
  • the random poset: for $(B,\leq)$ a finite poset, the map $b\mapsto\{b^{\prime}\in B:b^{\prime}\leq b\}$ is an embedding $(B,\leq)\to(\mathcal{P}(B),\subseteq)$. A straightforward application of the Hales-Jewett theorem for the alphabet $\{0,1\}$ and $|B|$-parameter sets gives point-Ramsey (see, for example, [29, Proposition 4.2.5]).
@@ 6.2. Equivalence relations
    If $i=\omega$, $j<\omega$, then $M_{i,j}$ has the externally definable Ramsey property but does not have point-Ramsey.
@@
    If $i<\omega,j=\omega$, then $M_{i,j}$ does not have the externally definable Ramsey property (and hence does not have point-Ramsey).
@@
    If $i=j=\omega$, then $M_{i,j}$ has point-Ramsey, and hence the externally definable Ramsey property.
@@
(1): by Proposition 3.13, we need only check for $|\bar{a}|=1$ and $\phi(x,y)\equiv x\sim y$. Let $\chi:\binom{M_{i,j}}{\bar{a}}\to 2$ be externally defined by some $e\in N\succeq M_{i,j}$ and $\phi(x,y)$. As all equivalence classes in $N$ are of size $\leq j$, we have $|\chi^{-1}(0)|\leq j$, and so using the EP of $M_{i,j}$, for $B\in\operatorname{Age}(M_{i,j})$ we find a copy $B^{\prime}\subseteq M_{i,j}$ of $B$ of colour $1$ in $\chi$, proving the claim.
To see that $M_{i,j}$ does not have vertex-Ramsey: take a $j$-colouring of the points of $M_{i,j}$ where for each equivalence class $E$ no two of the $j$ points of $E$ have the same colour.
@@
(3): It is straightforward to see that $M_{i,j}$ is indivisible, and hence by Lemma 2.18 and Lemma 2.19 we see that $M_{i,j}$ has point-Ramsey. ∎

### v4

@@ 3.2. Simplifications via quantifier elimination
Let $M$ be an ultrahomogeneous structure. Suppose $\operatorname{Th}(M)$ has QE. Let $\bar{a}\in\operatorname{Age}(M)$. Suppose $M$ has the $\bar{a}$-RP for externally atomic-definable $2$-colourings. Then M has the $\bar{a}$-RP for all externally definable $2$-colourings.
@@
Apply Lemma 3.8 to the set of atomic formulae. ∎
@@
Assume $M$ is ultrahomogeneous. We consider the case of equality. The $2$-colourings externally defined by the atomic formula $\phi(\bar{x},\bar{y})\equiv x_{0}=x_{1}$ or by $\phi(\bar{x},\bar{y})\equiv y_{0}=y_{1}$ are constant, so immediately we have that $M$ has the $\bar{a}$-RP for these colourings. However, it is possible for $M$ to not have the point-Ramsey property for $2$-colourings externally defined by the atomic formula $\phi(\bar{x},\bar{y})\equiv x=y$. For example, let $\mathcal{L}$ consist of a single unary predicate $R$, which we think of as the “red points” of a set, and let $M$ be the Fraïssé limit of the class of finite $\mathcal{L}$-structures $D$ with $|R^{D}|\leq 2$, i.e. with at most two red points. Let $a\in M$ be a red point, and let $B\in\operatorname{Age}(M)$ consist of two red points. Then the $2$-colouring of $\binom{M}{a}$ externally defined by $a\in M$ and the formula $\phi(x,y)\equiv x=y$ has no monochromatic copy of $B$.
However, if $\operatorname{Age}(M)$ has the strong JEP (the JEP with the additional condition that the images of the two embeddings are disjoint), then $M$ has the point-Ramsey property for $2$-colourings externally defined by $e\in M$ and $x=y$: any copy of $B$ which does not contain $e$ is monochromatic. Thus, if $\operatorname{Age}(M)$ has the strong JEP, $M$ automatically has the $\bar{a}$-RP for all $2$-colourings externally defined by atomic formulae involving equality.
@@
Let $M$ be an ultrahomogeneous $\mathcal{L}$-structure, where $\mathcal{L}$ has maximum arity $t\geq 2$. Suppose $\operatorname{Th}(M)$ has QE. Suppose $M$ has the externally definable $\bar{a}$-Ramsey property for all $\bar{a}\in\operatorname{Age}(M)$ with $|\bar{a}|<t$. Then $M$ has the externally definable Ramsey property.
@@
By Lemma 3.7 and Lemma 3.10, it suffices to consider externally atomic-definable $2$-colourings. Let $\bar{a}\in\operatorname{Age}(M)$, and let $\chi:\binom{M}{\bar{a}}\to 2$ be externally defined by some $\bar{e}\in N\succeq M$ and an atomic formula $\phi(\bar{x},\bar{y})$. Take subsequences $(x_{r(i)}),(y_{s(i)})$ of $\bar{x},\bar{y}$ respectively, consisting of the free variables that actually occur in $\phi(\bar{x},\bar{y})$. Let $u$ be the length of the subsequence $(x_{r(i)})$, and let $v$ be the length of the subsequence $(y_{s(i)})$.
@@
Let $M$ be an ultrahomogeneous $\mathcal{L}$-structure, where $\mathcal{L}$ has maximum arity $t\geq 2$. Suppose $\operatorname{Th}(M)$ has QE. Suppose, for $\bar{a}\in\operatorname{Age}(M)$ with $|\bar{a}|<t$, that $M$ has the $\bar{a}$-RP for externally atomic-definable $2$-colourings. Then $M$ has the externally definable Ramsey property.
(This resembles [15, Proposition 3.17].)
@@ 4. The fixed points on type spaces property (FPT)
Let $M$ be an $\mathcal{L}$-structure, and let $G=\operatorname{Aut}(M)$ with the pointwise convergence topology. Then $G$ acts continuously on the Stone spaces $S_{n}(M)$, with the action given by
@@
Note that we define the action of $G$ on $\mathcal{L}(M)$-formulae as $g\cdot\phi(\bar{x},\bar{m})=\phi(\bar{x},g\bar{m})$.
@@
Let $M$ be an $\mathcal{L}$-structure with automorphism group $G=\operatorname{Aut}(M)$. We say that $M$ has the fixed points on type spaces property (FPT) if every subflow of $G\curvearrowright S_{n}(M)$, $n\geq 1$, has a fixed point.
We also define a more restricted version of this property for particular $n$. Let $n\geq 1$. We say that $M$ has FPTn if every subflow of $G\curvearrowright S_{n}(M)$ has a fixed point.
@@ 5. FPT and the externally definable Ramsey property
    if $M$ has FPT, then $M$ has the externally definable Ramsey property;
@@
    if $\operatorname{Age}(M)$ is countable, then $M$ has the externally definable Ramsey property iff $M$ has FPT.
@@
(1): Let $\bar{a}\in\operatorname{Age}(M)$. We may assume $\bar{a}\subseteq M$. By Lemma 3.7, it suffices to show the $\bar{a}$-RP for externally definable $2$-colourings. Let $\chi:\binom{M}{\bar{a}}\to 2$ be externally defined by $\bar{e}\in N\succeq M$ and $\phi(\bar{x},\bar{y})$, where $\phi(\bar{x},\bar{y})$ has free variables among $\bar{x},\bar{y}$, but including all of $\bar{x}$. We show that for $B\in\operatorname{Age}(M)$, there exists a $\chi$-monochromatic copy $B^{\prime}\subseteq M$ of $B$.
Let $p(\bar{x})=\operatorname{tp}_{N}(\bar{e}/M)$. As $M$ has FPT, there is a $G$-invariant type $q(\bar{x})\in\overline{Gp(\bar{x})}$. Thus:
@@
(2): Take $p(\bar{x})\in S_{n}(M)$. We will show that there is a $G$-invariant type $q(\bar{x})\in\overline{Gp(\bar{x})}$.
Let $N\succeq M$ and $\bar{e}\in N$ be such that $\operatorname{tp}_{N}(\bar{e}/M)=p(\bar{x})$. Throughout this proof, when we specify externally defined colourings, the variables $\bar{x}$ will correspond to $\bar{e}$.
First, note that for any $\mathcal{L}$-formula $\phi(\bar{z})$, we may partition $\bar{z}$ into $\bar{z}\cap\bar{x}$ and $\bar{y}=\bar{z}\setminus\bar{x}$, and thereby consider $\phi(\bar{z})$ as a formula $\phi(\bar{x},\bar{y})$. For each $\mathcal{L}$-formula $\phi(\bar{x},\bar{y})$ and each $\bar{a}\in\operatorname{Age}(M)$ with $|\bar{a}|=|\bar{y}|$, let $\chi_{\phi,\bar{a}}:\binom{M}{\bar{a}}\to 2$ denote the colouring externally defined by $\bar{e}\in N\succeq M$ and $\phi(\bar{x},\bar{y})$. Note that for $\bar{a},\bar{b}\in\operatorname{Age}(M)$ with $\bar{a}\cong\bar{b}$, $\chi_{\phi,\bar{a}}=\chi_{\phi,\bar{b}}$.
@@
Let ${(\lambda_{r})}_{r<\omega}$ be given by the above Claim 2. Define $q(\bar{x})$ as follows. For $r<\omega$, let
$\psi_{r}(\bar{x},\bar{y})=\begin{cases}\phi_{r}(\bar{x},\bar{y})&\lambda_{r}=0\\ \neg\phi_{r}(\bar{x},\bar{y})&\lambda_{r}=1\end{cases}$
and add $\psi_{r}(\bar{x},\bar{a}^{\prime}_{r})$ to $q(\bar{x})$ for all $\bar{a}^{\prime}_{r}\in\binom{M}{\bar{a}_{r}}$. It is immediate that $q(\bar{x})$ is $G$-invariant. To show that $q(\bar{x})$ is a complete type, it suffices to show that $q(\bar{x})\cup\operatorname{Th}_{\mathcal{L}(M)}(M)$ is consistent. Take $s,t\in\mathbb{N}$. For $i<s,j<t$, take $\bar{a}_{i,j}\in\binom{M}{\bar{a}_{i}}$. We show that
$\{\psi_{i}(\bar{x},\bar{a}_{i,j}):i<s,j<t\}\cup\operatorname{Th}_{\mathcal{L}(M)}(M)$
is consistent, and this suffices by compactness. Take finite $B\subseteq M$ with $\bar{a}_{i,j}\subseteq B$ for $i<s,j<t$. Then, by Claim 2, there is $B^{\prime}\subseteq M$ and an isomorphism $f:B\to B^{\prime}$ with $\chi_{i}|_{B^{\prime}}=\{\lambda_{i}\}\text{ for }i<s$. Using ultrahomogeneity of $M$, there is $g\in\operatorname{Aut}(M)$ extending $f$. As $\chi_{i}|_{B^{\prime}}=\{\lambda_{i}\}$ for $i<s$, we have $N\models\psi_{i}(\bar{e},\bar{a}^{\prime}_{i})\text{ for }i<s\text{ and }\bar{a}^{\prime}_{i}\in\binom{B^{\prime}}{\bar{a}_{i}}$, and hence $N\models\psi_{i}(\bar{e},g\bar{a}_{i,j})\text{ for }i<s,j<t$, which gives the desired consistency.
@@
We note that, for the case of $\aleph_{0}$-saturated Fraïssé structures, Theorem 5.1 can be deduced from [15, Theorem 3.14] (personal communication, K. Krupiński). The proof of Theorem 5.1 given here uses elementary techniques, and applies to Fraïssé structures generally (or indeed ultrahomogeneous structures with countable age). Several of the examples in the following section are not $\aleph_{0}$-saturated, so [15, Theorem 3.14] cannot be applied, and here Theorem 5.1 becomes relevant. In Section 7.1, we give a general recipe for constructing such examples.
@@ 6. Examples
$\bar{a}$-Ramsey property for $|\bar{a}|<t$, where $t$ is the maximum arity of $\mathcal{L}$;
@@
6.1. Finite binary languages: the point-Ramsey property
@@
Let $\mathcal{L}$ be a language consisting of finitely many binary relation symbols. Let $M$ be a Fraïssé $\mathcal{L}$-structure with the point-Ramsey property. Then $M$ has the externally definable Ramsey property.
@@
By Lemma 2.6, $\operatorname{Th}(M)$ has QE. Thus, by Proposition 3.13, we need only check the point-Ramsey property for $2$-colourings of the points of $M$ which are externally defined by atomic formulae, and these form a subset of all $2$-colourings of points of $M$. As $M$ has the point-Ramsey property, we therefore have the externally definable Ramsey property. ∎
@@
The random graph has the point-Ramsey property.
@@
Denote the random graph by $M$. We show that $M$ has the pigeonhole property (see Definition 2.17), a stronger property than the point-Ramsey property. The following is from [3, Proposition 1.3.3] – we include it due to its brevity. Recall that the random graph has the witness property:
@@
One can also quickly prove the point-Ramsey property for the random graph via the lexicographic product: see [28, Theorem 12.1].
There are many other Fraïssé structures in finite binary relational languages with the point-Ramsey property, and therefore the externally definable Ramsey property. For example:
  • the random tournament: this structure has the pigeonhole property by an argument as for the random graph, and we can also prove the point-Ramsey property via the lexicographic product;
  • the $K_{n}$-free random graph ($n\geq 3$): this follows via Folkman’s theorem, the proof of which adapts the lexicographic product proof for the random graph ([7]; a clear exposition following [14] is [28, Theorem 12.3]); alternatively, we may use the Nešetřil-Rödl theorem as in Section 6.3;
  • the random poset: for $(B,\leq)$ a finite poset, the map $b\mapsto\{b^{\prime}\in B:b^{\prime}\leq b\}$ is an embedding $(B,\leq)\to(\mathcal{P}(B),\subseteq)$. A straightforward application of the Hales-Jewett theorem for the alphabet $\{0,1\}$ and $|B|$-parameter sets gives the point-Ramsey property (see, for example, [28, Proposition 4.2.5]).
@@ 6.2. Equivalence relations
    If $i=\omega$, $j<\omega$, then $M_{i,j}$ has the externally definable Ramsey property but does not have the point-Ramsey property.
@@
    If $i<\omega,j=\omega$, then $M_{i,j}$ does not have the externally definable Ramsey property (and hence does not have the point-Ramsey property).
@@
    If $i=j=\omega$, then $M_{i,j}$ has the point-Ramsey property, and hence the externally definable Ramsey property.
@@
(1): By Proposition 3.13 and Remark 3.11 (note that $\operatorname{Age}(M_{i,j})$ has the strong JEP), we need only check the externally definable RP for $|\bar{a}|=1$ and $\phi(x,y)\equiv x\sim y$. Let $\chi:\binom{M_{i,j}}{\bar{a}}\to 2$ be externally defined by some $e\in N\succeq M_{i,j}$ and $\phi(x,y)$. As all equivalence classes in $N$ are of size $\leq j$, we have $|\chi^{-1}(0)|\leq j$, and so using the EP of $M_{i,j}$, for $B\in\operatorname{Age}(M_{i,j})$ we find a copy $B^{\prime}\subseteq M_{i,j}$ of $B$ of colour $1$ in $\chi$, and hence $M_{i,j}$ has the externally definable RP.
To see that $M_{i,j}$ does not have the point-Ramsey property: take a $j$-colouring of the points of $M_{i,j}$ where for each equivalence class $E$ no two of the $j$ points of $E$ have the same colour.
@@
(3): It is straightforward to see that $M_{i,j}$ is indivisible, and hence by Lemma 2.18 and Lemma 2.19 we see that $M_{i,j}$ has the point-Ramsey property. ∎